OpenStudy (anonymous):

Derek kicks a soccer ball off the ground and in the air with an initial velocity of 31 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches?

OpenStudy (cwrw238):

i can do this by using v = u + at u = 31 ft/sec, v = 0 (at maximum height, ball is stationary), a = -32 ft / sec/sec 0 = 31 - 32 t t = 31/32 seconds height = u + v ----- * t 2 = 31 31 -- * -- = 15.02 ft 2 32

OpenStudy (anonymous):

Thank you!! :)

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

i'm confused by your formula - what is s?